# Graph Neural Preconditioners for Iterative Solutions of Sparse Linear Systems

### v1

@@ Appendix B Additional Experiment Results
In Figure 7, we plot the proportion of problems for which each preconditioner performs the best, for different matrix sizes and condition numbers. This figure is a counterpart of Figure 3. The main observation is that the proportion of problems for which GNP performs the best is relatively flat across matrix sizes. Meanwhile, the right plot shows that GNP is particularly useful for ill-conditioned problems.
@@
Figure 7: Proportion of problems for which each preconditioner performs the best, for different matrix sizes and condition numbers. The last bar on the right plot is for condition number $\geq 10^{16}$. Note that not every matrix has a known condition number.
@@
In Figure 8, we show the convergence histories of the systems appearing in Figure 4, with respect to both iteration counts and time. We also plot the training curves. The short solution time of GNP is notable, especially when compared with the solution time of GMRES as a preconditioner. The training loss generally stays on the level of 1e-2 to 1e-3. For some problems (e.g., VanVelzen_std1_Jac3), training does not seem to progress well, but the training loss is already at a satisfactory point in the first place, and hence the preconditioner can be rather effective. We speculate that this behavior is caused by the favorable architectural inductive bias in the GNN.

### v3

@@ F.3 Breakdown of Best Preconditioners With Respect to Matrix Distributions
In Figure 9, we plot the proportion of problems on which each preconditioner performs the best, for different matrix sizes, condition numbers, and application areas. This figure is a counterpart of Figure 4. The main observation is that the proportion of problems on which GNP performs the best is relatively flat across matrix sizes. Other observations follow closely from those of Figure 4; namely, GNP is particularly useful for ill-conditioned problems and some application areas (chemical process simulation problems, economic problems, and eigenvalue/model reduction problems).
@@
Figure 9: Proportion of problems on which each preconditioner performs the best, for different matrix sizes, condition numbers, and application areas. Only the application areas with the top number of problems are shown. The last bar in the middle plot is for condition number $\geq 10^{16}$. Note that not every matrix has a known condition number.
F.4 Comparison between PyAMG and AmgX
Besides PyAMG, another notable implementation of AMG (which supports using it as a preconditioner and supports GPU) is AmgX https://github.com/NVIDIA/AMGX. Here, we compare the two implementations in terms of robustness and speed.
Unlike PyAMG, which is a blackbox, AmgX leaves the choice of the AMG method and the parameters to the user. We used the driver examples/amgx_capi.c and the configuration src/configs/FGMRES_AGGREGATION.json, which appears to be the most reasonable for the diverse matrices and application areas at hand. This setting uses aggregation-based AMG rather than classical AMG, which is also consistent with the preference expressed by PyAMG.
Here are some findings.
AmgX is less robust than PyAMG. Table 3 shows that AmgX causes more failures during the construction as well as the solution phase. In fact, there were three cases where AmgX hanged, which hampered automated benchmarking.
Table 3: Failures of two implementations of AMG.
PyAMG AmgX
Construction failure 62 (7.15%) 105 (12.11%)
Solution failure 5 (0.58%) 36 (4.15%)
Despite a lack of robustness, the construction of AmgX is much faster than that of PyAMG. For AmgX, the distribution of the preconditioner construction time has a minimum 0.006s, median 0.017s, and maximum 4.895s. In contrast, the distribution for PyAMG has a minimum 0.005s, median 0.286, and maximum 8505.1s. Nevertheless, there are still 16% of the cases where the PyAMG preconditioner is faster to construct than the AmgX preconditioner.
We skip the comparison of the solution time. This is because in AmgX, the Krylov solver and the preconditioner are bundled inside a C complementation. A fair comparison requires isolating the preconditioner from the solver, because our benchmarking framework implements the solver in Python. However, the isolation is beyond scope due to the complexity of the implementation, such as the use of data structures and cuda handling. Nevertheless, we expect that the solution time of AmgX is faster than that of PyAMG, based on the observations made for preconditioner construction.
F.5 Changing the Default PyAMG Solver
The blackbox solver of PyAMG, pyamg.blackbox.solver, uses the classical-style smoothed aggregation method pyamg.aggregation.smoothed_aggregation_solver (SA). For nonsymmetric matrices, however, the AIR method works better (Manteuffel et al., 2019; 2018). In this experiment, we modify the blackbox solver by calling pyamg.classical.air_solver for nonsymmetric matrices.
We first tried using the default options of AIR, but encountered “zero diagonal encountered in Jacobi” errors. These errors were caused by the use of jc_jacobi as the post-smoother. Then, we replaced the smoothers with the default ones used by SA (gauss_seidel_nr). In this case, the solver hung on the problem FIDAP/ex40. Before this, 105 problems were solved, among which 83 were nonsymmetric. The results of these 83 systems are summarized in Table 4, suggesting that AIR is indeed better but not always.
Table 4: Counts of problems when comparing AIR with SA for nonsymmetric matrices.
AIR is is better than SA AIR is worse The two are equal
By iteration count 14 8 61
By final residual 50 29 4
Overall, we can conclude that AIR has not exhibited its full potential for nonsymmetric matrices. We speculate that a robust implementation of AIR is challenging and that might be the reason why the authors did not use it for the blackbox solver in the first place.
F.6 Convergence Histories
@@
In Figure 10, we show the convergence histories of the systems appearing in Figure 5, with respect to both iteration count and time. We also plot the training curves. The short solution time of GNP is notable, especially when compared with the solution time of GMRES as a preconditioner. The training loss generally stays on the level of 1e-2 to 1e-3. For some problems (e.g., VanVelzen_std1_Jac3), training does not seem to progress well, but the training loss is already at a satisfactory point in the first place, and hence the preconditioner can be rather effective. We speculate that this behavior is caused by the favorable architectural inductive bias in the GNN.
@@
To provide further examples, we show the convergence histories of all systems in the application areas identified by Figure 4 to favor GNP: chemical process simulation problems (Figures 1113), economic problems (Figures 1416), and eigenvalue/model reduction problems (Figures 1718). In many occasions, when GNP performs the best, it outperforms the second best substantially. It can even decrease the relative residual norm under rtol while other preconditioners barely solve the system.
Figure 11: Convergence of the linear system solutions for chemical process simulation problems (1/3).
Figure 12: Convergence of the linear system solutions for chemical process simulation problems (2/3).
Figure 13: Convergence of the linear system solutions for chemical process simulation problems (3/3).
Figure 14: Convergence of the linear system solutions for economic problems (1/3).
Figure 15: Convergence of the linear system solutions for economic problems (2/3).
Figure 16: Convergence of the linear system solutions for economic problems (3/3).
Figure 17: Convergence of the linear system solutions for eigenvalue/model reduction problems (1/2).
Figure 18: Convergence of the linear system solutions for eigenvalue/model reduction problems (2/2).